# SS Linear Fusion Model: Hyperspectral Imaging Efficient Spatial and Spectral Linear Model with Bidirectional Feature Learning

### v1

@@ IV-D3 Efficiency Verification
Efficiency is critical in hyperspectral image classification, where large data volumes can lead to high computational demands. In this section, we assess the SS Linear Model’s efficiency across three key metrics: memory consumption, training and testing durations, and overall computational complexity. Our analysis uses the Houston 2013 dataset with a standardized input size of 1×15×15×200, allowing for a fair comparison with other leading models. Table XI provides a summary of the computational requirements across competing methods.
@@
Floating Point Operations (FLOPs): The SS Linear Model exhibits significantly lower FLOPs (0.05 GB), the smallest of all models tested. This low FLOP count results from the model’s streamlined bidirectional processing, which avoids complex self-attention mechanisms. Compared to transformer-based methods such as T2T and Deep ViT, which require 5.95 GB and 2.71 GB, respectively, the SS Linear Model demonstrates a marked reduction in computational demand. This makes it ideal for deployment in environments with limited processing power, such as edge devices or mobile platforms.
Parameter Count and Memory Efficiency: While maintaining high accuracy, the SS Linear Model has a compact parameter count (50.32 MB), comparable to 3D-CNN (1.46 MB) but considerably more efficient than T2T (730.18 MB). This smaller parameter set enables faster access and reduced memory usage, an essential factor for large-scale hyperspectral analysis where memory constraints often limit model choices.
Training and Testing Time: The SS Linear Model achieves a training time of 130.30 seconds and a testing time of 1.90 seconds, positioning it as one of the most efficient models in terms of both metrics. This efficiency is particularly relevant for real-time applications, where rapid model updates and quick inference times are essential. Although transformer models like T2T and HiT provide strong performance, they incur significantly higher training and testing times (e.g., T2T requires 479.91 seconds for training and 6.79 seconds for testing).
Comparison with CNN-Based Models: CNN-based methods such as 2D-CNN and R-2D-CNN show competitive training and testing times; however, their lower computational complexity often comes at the expense of reduced accuracy. In contrast, the SS Linear Model balances low computational complexity with high accuracy, surpassing traditional CNNs in classification performance while maintaining similar efficiency levels.
Practical Implications for Deployment: The computational efficiency of SS Linear Model, particularly Method 2 with patch size 5, allows it to maintain high accuracy with reduced patch size while minimizing computational demands. This positions the model as an optimal choice for applications that require rapid, accurate analysis of hyperspectral data, such as precision agriculture, real-time urban monitoring, and mobile-based remote sensing tasks.
The efficiency verification underscores SS Linear Model’s suitability for real-world deployment, where both high accuracy and computational efficiency are essential. The model’s low FLOPs, compact memory footprint, and fast processing times set it apart from traditional CNN and transformer models, making it a powerful yet resource-conscious option for hyperspectral image classification.
@@ V Conclusions
This study introduced the SS Linear Model, a bidirectional state-space model designed specifically for hyperspectral image classification. By integrating bidirectional state-space processing and efficient spectral-spatial feature extraction, the SS Linear Model addresses two primary challenges in hyperspectral imaging: the high-dimensional nature of spectral data and the computational complexities of both spectral and spatial information processing. This model not only compresses hyperspectral data efficiently, but also enhances classification performance through its unique architecture, setting a new benchmark for hyperspectral analysis.
@@
Comprehensive evaluations on three major hyperspectral datasets: Houston, India, and the University of Pavia demonstrate that the SS Linear Model consistently outperforms conventional transformer-based and CNN-based approaches. It achieves high accuracy with reduced computational demand, underscoring its robustness and adaptability across diverse landscapes and land cover types. This superior classification performance is achieved without the resource intensity typically associated with advanced transformer models, showcasing the model’s efficiency and accuracy trade-offs.
@@
A significant feature of the SS Linear Model is its streamlined computational efficiency. Through careful design that minimizes memory requirements and computational operations, the model achieves rapid training and testing, making it suitable for deployment in environments with limited computing power. Experimental results indicate that the SS Linear Model achieves similar or improved accuracy with substantially lower FLOPs and parameter counts, thereby opening possibilities for large-scale, real-time hyperspectral data analysis in resource-constrained settings.
@@
This efficiency positions the SS Linear Model as a practical choice for applications requiring high-speed, accurate hyperspectral analysis, such as precision agriculture, urban monitoring, and mobile-based remote sensing. By achieving a balance between accuracy and computational demand, the SS Linear Model represents a critical advancement in making sophisticated hyperspectral imaging technology accessible and operationally feasible across various platforms.
@@
In conclusion, the SS Linear Model stands as a pioneering model that combines bidirectional processing with computational efficiency, advancing the field of hyperspectral imaging. Its design exemplifies the potential of state-space models for high-dimensional spectral data analysis, paving the way for future research into their broader application in remote sensing. This model not only advances hyperspectral imaging capabilities but also inspires further investigation into the use of state-space models for scalable, efficient, and accurate remote sensing data analysis. The results of this study support the role of the SS linear model as a transformative approach, promising continued exploration of its utility and impact across the spectrum of remote sensing technologies.

### v2

@@ IV-D3 Efficiency Verification
Efficiency is critical in hyperspectral image classification, where large data volumes can lead to high computational demands. In this section, we assess the SS non-Linear Model’s efficiency across three key metrics: memory consumption, training and testing durations, and overall computational complexity. Our analysis uses the Houston 2013 dataset with a standardized input size of 1×15×15×200, allowing for a fair comparison with other leading models. Table XI provides a summary of the computational requirements across competing methods.
@@
Floating Point Operations (FLOPs): The SS non-Linear Model exhibits significantly lower FLOPs (0.05 GB), the smallest of all models tested. This low FLOP count results from the model’s streamlined bidirectional processing, which avoids complex self-attention mechanisms. Compared to transformer-based methods such as T2T and Deep ViT, which require 5.95 GB and 2.71 GB, respectively, the SS non-Linear Model demonstrates a marked reduction in computational demand. This makes it ideal for deployment in environments with limited processing power, such as edge devices or mobile platforms.
Parameter Count and Memory Efficiency: While maintaining high accuracy, the SS non-Linear Model has a compact parameter count (50.32 MB), comparable to 3D-CNN (1.46 MB) but considerably more efficient than T2T (730.18 MB). This smaller parameter set enables faster access and reduced memory usage, an essential factor for large-scale hyperspectral analysis where memory constraints often limit model choices.
Training and Testing Time: The SS non-Linear Model achieves a training time of 130.30 seconds and a testing time of 1.90 seconds, positioning it as one of the most efficient models in terms of both metrics. This efficiency is particularly relevant for real-time applications, where rapid model updates and quick inference times are essential. Although transformer models like T2T and HiT provide strong performance, they incur significantly higher training and testing times (e.g., T2T requires 479.91 seconds for training and 6.79 seconds for testing).
Comparison with CNN-Based Models: CNN-based methods such as 2D-CNN and R-2D-CNN show competitive training and testing times; however, their lower computational complexity often comes at the expense of reduced accuracy. In contrast, the SS non-Linear Model balances low computational complexity with high accuracy, surpassing traditional CNNs in classification performance while maintaining similar efficiency levels.
Practical Implications for Deployment: The computational efficiency of SS non-Linear Model, particularly Method 2 with patch size 5, allows it to maintain high accuracy with reduced patch size while minimizing computational demands. This positions the model as an optimal choice for applications that require rapid, accurate analysis of hyperspectral data, such as precision agriculture, real-time urban monitoring, and mobile-based remote sensing tasks.
The efficiency verification underscores SS non-Linear Model’s suitability for real-world deployment, where both high accuracy and computational efficiency are essential. The model’s low FLOPs, compact memory footprint, and fast processing times set it apart from traditional CNN and transformer models, making it a powerful yet resource-conscious option for hyperspectral image classification.
@@ V Conclusions
This study introduced the SS non-Linear Model, a bidirectional state-space model designed specifically for hyperspectral image classification. By integrating bidirectional state-space processing and efficient spectral-spatial feature extraction, the SS non-Linear Model addresses two primary challenges in hyperspectral imaging: the high-dimensional nature of spectral data and the computational complexities of both spectral and spatial information processing. This model not only compresses hyperspectral data efficiently, but also enhances classification performance through its unique architecture, setting a new benchmark for hyperspectral analysis.
@@
Comprehensive evaluations on three major hyperspectral datasets: Houston, India, and the University of Pavia demonstrate that the SS non-Linear Model consistently outperforms conventional transformer-based and CNN-based approaches. It achieves high accuracy with reduced computational demand, underscoring its robustness and adaptability across diverse landscapes and land cover types. This superior classification performance is achieved without the resource intensity typically associated with advanced transformer models, showcasing the model’s efficiency and accuracy trade-offs.
@@
A significant feature of the SS non-Linear Model is its streamlined computational efficiency. Through careful design that minimizes memory requirements and computational operations, the model achieves rapid training and testing, making it suitable for deployment in environments with limited computing power. Experimental results indicate that the SS non-Linear Model achieves similar or improved accuracy with substantially lower FLOPs and parameter counts, thereby opening possibilities for large-scale, real-time hyperspectral data analysis in resource-constrained settings.
@@
This efficiency positions the SS non-Linear Model as a practical choice for applications requiring high-speed, accurate hyperspectral analysis, such as precision agriculture, urban monitoring, and mobile-based remote sensing. By achieving a balance between accuracy and computational demand, the SS non-Linear Model represents a critical advancement in making sophisticated hyperspectral imaging technology accessible and operationally feasible across various platforms.
@@
In conclusion, the SS non-Linear Model stands as a pioneering model that combines bidirectional processing with computational efficiency, advancing the field of hyperspectral imaging. Its design exemplifies the potential of state-space models for high-dimensional spectral data analysis, paving the way for future research into their broader application in remote sensing. This model not only advances hyperspectral imaging capabilities but also inspires further investigation into the use of state-space models for scalable, efficient, and accurate remote sensing data analysis. The results of this study support the role of the SS non-Linear Model as a transformative approach, promising continued exploration of its utility and impact across the spectrum of remote sensing technologies.